OpenStudy (anonymous):

Every morning, Jorge runs a circular path that has a radius of 42 ft. How far does Jorge run if he does one lap around the track?

OpenStudy (anonymous):

the distance covered in these cases equals the circumference hence reqd distance= 2*22/7*42 ft=2*22*6=264 ft

OpenStudy (anonymous):

Thanks you're a life saver!